# Meiying Zhang1, Weiyuan Peng1, Guangyao Ding2, Chenyang Lei2, Chunlin Ji3, Qi Hao2

### v1

@@ I Introduction
  • Geneartion of High-quality Pseudo-label. The object size bias and distribution differences between the simulated and real data, as shown in Fig 1, easily lead to inconsistent regression results (i.e., low-quality pseudo-labels). How to mitigate these biases in detection is important for generating high-quality pseudo-labels.
@@
To reduce the domain gap arising from object bias, current methods primarily focus on point cloud preprocessing in the source domain. However, these methods can barely reduce domain inconsistencies between two domains [13, 8, 7]. Furthermore, methods that use a complex two-stage UDA design show limited performance in sim-to-real tasks [13, 6]. Meanwhile, various methods have been proposed to achieve high-quality pseudo-label guidance, including multi-output fusion techniques, such as fusing multi-modality outputs for 2D-3D data [12], or fusing multi-pass outputs to maintain “high stochasticity” [14]. The mean teacher scheme can also generate more accurate pseudo-labels in target domains [14, 6, 15]. However, its performance can be much degraded by the data noise in sim-to-real tasks.
@@
  • Development of a uniform corner-format measure for aleatoric uncertainty (AU) estimation to evaluate the quality pseudo-labels accurately.
  • Development of two sampling strategies based on AU in the mean teacher domain adaptation process to choose those point cloud frames and labels with adequate label quality only.
  • Release of the open source code of CTS, alongside the CARLA3D simulated dataset, for further research [^1] The code of CTS and CARALA3D dataset are available at https://github.com/tendo518/CTS-UDA.
@@ II-A UDA for 3D object detection
Some previous works have well explored the usage of UDA in 3D object detection [7, 8, 16, 13, 6, 14]. One common challenge of UDA in 3D object detection is the object size bias when cross-domains. Wang et al.[13] propose statistical normalization (SN) to align object sizes utilizing statistical information from target domain data. ST3D[7] and ST3D++[8] employ data augmentations during source domain training to improve the model’s incorporation of diverse size information. Besides mitigating object size bias, using pseudo-label guided methods in UDA emphasizes improving the quality of pseudo-labels. JST [12] enhances pseudo-label quality through 2D and 3D joint refinement, aligning outcomes from both modalities. ST3D [7] integrates an additional IoU regression head to assess prediction quality, facilitating selective updates of the pseudo-label pool. Building upon ST3D, ST3D++ [8] further refines pseudo-labels using a quality-aware denoising pipeline. MLC-Net[6] also employs the mean teacher scheme to ensure target domain consistency between teacher and student modules at both point and instance levels, which is similar to our method but involves higher complexity using UDA design for both stages. Although having significant improvements in real-to-real tasks, existing UDA methods often experience serious performance degradation in sim-to-real tasks. Therefore, based on the analysis of simulation and reality differences, our study concentrates on the quality enhancement, evaluation and selection for pseudo-labels to achieve higher sim-to-real performance.
@@ II-B Uncertainty Estimation in 3D Object Detection
Uncertainty can serve as a valuable metric for quantifying both data and model noise within deep neural networks (DNNs) [17, 18, 19, 20, 21]. Uncertainty estimation methods typically address two main sources: epistemic uncertainty (EU) and aleatoric uncertainty (AU). EU is represented by a posterior distribution over model parameters[17, 18, 20], providing insights into the models uncertainty; AU is represented a distribution over model outputs[19, 21], reflecting intrinsic data stochasticity. Notably, AU varies with the quality of input data, suitable for quantifying the noise level of input data. Within the context of 3D detection tasks, several methodologies have integrated aleatoric uncertainty (AU) due to its ability to enhance detection performance[22, 23, 24, 25]. Meyer et al.[22] employ a mixture of Laplace distributions to fit the variances for each predefined regression variable, including box center positions, sizes, and orientation. Feng et al.[24] model AU using multivariate Gaussian distributions, with independent variables representing three distinct sets, i.e., RoI positions, bounding box positions, and orientation. However, few methods have leveraged the Aleatoric Uncertainty (AU) estimated from 3D detection results for the evaluation of data noise. Besides, existing approaches represent uncertainties using non-uniform variables, adding a complexity to further utilization. Therefore, this study proposes a uniform corner-based representation for bounding boxes with uncertainties, easy for the quality evaluation of the predicted psudo-labels.
@@ III System Setup
The complete CTS framework diagram is shown in Fig 2. The CTS framework leverages simulated data from the source domain to develop detection capabilities, followed by model refinement through mean teacher-based domain adaptation in real traffic scenarios of the target domain. The mean teacher scheme involves two branches: the student and teacher models. They share identical architectures and are both initialized with parameters from the source-domain training. However, they undergo different update mechanisms:
Student Model: The student model utilizes augmented RoI points and features as input, supervised with pseudo-labels $\hat{y}_{t}$ on the target domain or ground truth labels $y_{s}$ in the source domain. It is worth noting that the generated pseudo-labels can serve as supervision for the 1st-stage network as well, thus enabling domain adaptation for the 1st-stage network. Thus, the total loss of this network includes: 1. first-stage RoI regression loss $l_{reg1}$. 2. first-stage RoI classification loss $l_{cls1}$. 3. second-stage regression Smooth-L1 loss $l_{reg2}$. 4. second-stage classification loss $l_{cls2}$. 5. second-stage AU-NLL loss $l_{nll}$ specified in Sec IV-B.
Teacher Model: The teacher model handles raw (non-augmented) data and keeps its weights fixed. Instead of standard backpropagation, it updates its weights using the exponential moving average (EMA):
@@
where $\theta^{stu}_{t}$ is the student weight, $\beta$ is the EMA decay that controls the updating ratio and $t$ stand for $t_{th}$ iteration.
@@ IV-A1 Anchor Head (AH)
The second-stage model typically predicts size residuals $\Delta_{whl}$ between proposals from the first-stage and final bounding boxes, denoted as $\hat{B}$. This approach avoids regressing the size of bounding boxes entirely from scratch. However, a challenge arises when the first-stage model, trained with biased supervision from source domain labels, exhibits inaccuracy in estimating proposal sizes. Unreliable proposal box sizes can lead to size errors accumulating in the second stage, degrading final bounding box refinement accuracy and the effectiveness of pseudo-labels. Inspired by anchor-based detectors [26], we introduce a fixed-size anchor box ${w_{an},h_{an},l_{an}}$ to replace the proposal, termed the anchor head (AH). With AH, the second-stage network no longer refines proposals, but rather operates on a globally fixed-size 3D anchor. By employing AH in both the source and target domain training, AH ensures consistency in the second-stage network’s behavior across domains, thereby facilitating domain adaptation and improving the quality of pseudo-labels.
@@ IV-A2 RoI Random Scaling (RRS) and Augmentation
where $\theta$ is the model parameter, $f_{\mu}$ and $f_{\sigma^{2}}$ represent sub-networks for predicting the mean and the variance.
@@
$\mathcal{L}_{nll}^{i}=\frac{\left(\overline{\mathbf{y}_{c}^{i}}-\overline{\hat{\mathbf{y}}_{c}^{i}}\right)^{2}}{2(\sigma_{c}^{i})^{2}}+\frac{1}{2}\log(\sigma_{c}^{i})^{2}$ $\mathcal{L}_{nll}=\frac{1}{8}\sum_{i=1}^{8}\mathcal{L}_{nll}^{i}$ (5)
$u_{box}=\frac{1}{8}\sum_{i=1}^{8}(\sigma_{c}^{i})^{2}$ (6)
And all components contribute equally to the loss and final uncertainty metric.
$(x,y,z)$${h}$${w}$${l}$${\alpha}$
(a) Encoded with BF
32674158
(b) Encoded with CF
@@ IV-C Noise-aware Mean Teacher
Aligning transformations on both student-model inputs and teacher-model output facilitates the acquisition of domain-invariant representations, thereby aiding in adaptation to the target domain using pseudo-labels. However, noisy pseudo-labels can lead to error accumulation. To address this challenge, we leverage aleatoric uncertainties predicted by a model to annotate data in the target domain and mitigate the impact of noisy data during mean teacher domain adaptation with following sampling strategies:
@@ IV-C2 Frame-Level Sampling
1:$\mathcal{T}$: Unlabeled Target Domain Dataset
2:$\mathcal{T}_{sub}$: Target Sub-dataset after Sampling
3:$N_{t}$: Number of samples in $\mathcal{T}$
[…]
5:${U}_{frame}\leftarrow\{\}$, $\mathcal{T}_{sub}\leftarrow\{\}$Foreach frame $x^{t}$ in $\mathcal{T}$
6:$\hat{\mathbf{y}}^{t},\hat{\mathbf{u}}^{t}\leftarrow$ inference $\mathcal{D}$ for $x^{t}$
7:$\hat{u}^{t}\leftarrow\text{mean of }\hat{\mathbf{u}}^{t}\text{ for all valid object in $x^{t}$}$
[…]
9:$j\leftarrow\text{argmin of }{U}_{frame}$
10:$\mathcal{T}_{sub}\leftarrow$ append the $j_{th}$ element $x_{j}^{t}$ in $\mathcal{T}$ to $\mathcal{T}_{sub}$
11:pop the $j_{th}$ element $\hat{u}_{j}^{t}$ from ${U}_{frame}$EndFor
12:$\mathcal{D}\leftarrow$ fine-tune $\mathcal{D}$ with $\mathcal{T}_{sub}$
13:${N}_{sub}+={N}_{sub}$EndWhile
14:return $\mathcal{D}$
[…]
@@ V-A1 Datasets
We conduct supervised training in a simulated source domain, namely CARLA3D, acquired within the CARLA simulator [9]. All samples are taken from eight built-in scenarios in CARLA to ensure data diversity. The ego-vehicle is positioned randomly, collecting about 100 samples per scenario, each comprising eight frames at 2Hz. Out of the eight frames per sample, five are randomly chosen for the training set, yielding 3,990 frames with a total of 25,192 objects. Further details of the CARLA3D dataset are outlined in Table II. The target domains chosen include KITTI [10], Lyft [11], and TinySUScape used in [12]. During the testing phase, samples from these datasets along with their corresponding labels will be utilized, whereas only samples will be used during the training phase. A summary of these datasets is presented in Table III.
@@ V-B Main Results
Our CTS framework was compared with the following methods: 1. SN[13]: A domain adaptation method has been considered effective on various datasets; 2. MLC-Net[6]: A domain adaptation method also based on mean teacher, which is similar to ours in the mean teacher part; 3. ST3D++[8]: A recent self-training based method that achieved state-of-the-art performance in real-to-real (e.g., Nusenses[31] $\to$ KITTI[10]) domain adaptation tasks.
@@ V-C4 Benefits of Noise Awareness in Mean Teacher
Figure 6: Examples of different levels of difficulties in 3D object detection. The blue boxes represent the ground truth; the green boxes represent the predicted results. The points in different colors at the box corners represent the 8 AU value components, whose mean is the final AU value of the entire object.
@@ VI Conclusion
This paper has presented a CTS framework for unsupervised domain adaptation (UDA) in 3D object detection from simulation to real-world domains. The proposed techniques including RoI random scaling and augmentation as well as the fixed-size archor head can enrich the diversity of simulation data and mitigate object size deviation across domains, respectively, enhancing the quality of pseudo-labels. The proposed aleatoric uncertainty (AU) estimation based uniform corner-format representation of bounding boxes can help integrate the awareness of pseudo-label noise into the mean teacher domain adaptation process, achieving high-quality pseudo-label sampling. Experimental results using the CARLA, KITTI, Lyft, and TinySUScape datasets can demonstrate significant improvements over existing methods in various sim-to-real UDA tasks, including 5%-17% improvement in $AP_{3D}$ and 2%-10% improvement in $AP_{BEV}$. Our future work will focus on extending our method to encompass both sim-to-real and real-to-real UDA scenarios.

### v2

@@ I Introduction
  • Generation of High-quality Pseudo-label. The object size bias and distribution differences between the simulated and real data, as shown in Fig 1, easily lead to inconsistent regression results (i.e., low-quality pseudo-labels). How to mitigate these biases in detection is important for generating high-quality pseudo-labels.
@@
To reduce the domain gap arising from object bias, current methods primarily focus on point cloud preprocessing in the source domain. However, these methods can barely reduce domain inconsistencies between two domains [13, 8, 7]. Furthermore, methods that use a complex two-stage UDA design show limited performance in sim-to-real tasks [13, 6]. Meanwhile, various methods have been proposed to achieve high-quality pseudo-label guidance, including multi-output fusion techniques, such as fusing multi-modality outputs for 2D-3D data [12], or fusing multi-pass outputs to maintain “high stochastic” [14]. The mean teacher scheme can also generate more accurate pseudo-labels in target domains [14, 6, 15]. However, its performance can be much degraded by the data noise in sim-to-real tasks.
@@
  • Development of a uniform corner-format measure for aleatoric uncertainty (AU) estimation to evaluate the quality of pseudo-labels accurately.
  • Development of two AU-based sampling strategies in the mean teacher domain adaptation process to select point cloud frames and labels with adequate quality.
  • Release of CTS code, alongside the CARLA3D simulated dataset, for further research [^1] The code of CTS and CARALA3D dataset are available at https://github.com/tendo518/CTS-UDA.
@@ II-A UDA for 3D object detection
Some previous works have well explored the usage of UDA in 3D object detection [7, 8, 13, 6, 14]. One common challenge of UDA in 3D object detection is the object size bias when cross-domains. Wang et al.[13] propose statistical normalization (SN) to align object sizes utilizing statistical information from target domain data. ST3D[7] and ST3D++[8] employ data augmentations during source domain training to improve the model’s incorporation of diverse size information. Besides mitigating object size bias, using pseudo-label guided methods in UDA emphasizes improving the quality of pseudo-labels. JST [12] enhances pseudo-label quality through 2D and 3D joint refinement, aligning outcomes from both modalities. ST3D [7] integrates an additional IoU regression head to assess prediction quality, facilitating selective updates of the pseudo-label pool. Building upon ST3D, ST3D++ [8] further refines pseudo-labels using a quality-aware denoising pipeline. MLC-Net[6] also employs the mean teacher scheme to ensure target domain consistency between teacher and student modules at both point and instance levels, which is similar to our method but involves higher complexity using UDA design for both stages. Although having significant improvements in real-to-real tasks, existing UDA methods often experience serious performance degradation in sim-to-real tasks. Therefore, based on the analysis of simulation and reality differences, our study concentrates on the quality enhancement, evaluation and selection for pseudo-labels to achieve higher sim-to-real performance.
@@ II-B Uncertainty Estimation in 3D Object Detection
Uncertainty can serve as a valuable metric for quantifying both data and model noise within deep neural networks (DNNs) [16, 17, 18, 19, 20]. Uncertainty estimation methods typically address two main sources: epistemic uncertainty (EU) and aleatoric uncertainty (AU). EU is represented by a posterior distribution over model parameters[16, 17, 19], providing insights into the models’ uncertainty; AU is represented a distribution over model outputs[18, 20], reflecting intrinsic data stochastic. Notably, AU varies with the quality of input data, suitable for quantifying the noise level of input data. Within the context of 3D detection tasks, several methodologies have integrated aleatoric uncertainty (AU) due to its ability to enhance detection performance[21, 22]. Meyer et al.[21] employ a mixture of Laplace distributions to fit the variances for each predefined regression variable, including box center positions, sizes, and orientation. Feng et al.[22] model AU using multivariate Gaussian distributions, with independent variables representing three distinct sets, i.e., RoI positions, bounding box positions, and orientation. However, few methods have leveraged the Aleatoric Uncertainty (AU) estimated from 3D detection results for the evaluation of data noise. Besides, existing approaches represent uncertainties using non-uniform variables, adding a complexity to further utilization. Therefore, this study proposes a uniform corner-based representation for bounding boxes with uncertainties, easy for the quality evaluation of the predicted pseudo-labels.
@@ III System Setup
The complete CTS framework is illustrated in Fig. 2. This framework utilizes simulated data from the source domain to initially develop detection capabilities, followed by model refinement through mean teacher-based domain adaptation in real-world traffic scenarios of the target domain. The mean-teacher method [5] mitigates the impact of domain shift by training the student model with a consistency objective, effectively utilizing unlabeled data from the target domain to improve the model’s performance in that domain.
The framework consists of two branches: the student and teacher models. Both models share the same architecture and are initialized with parameters trained on the source domain. However, they are updated through different mechanisms:
III-1 Student Model
The student model utilizes augmented RoI points and features as input, supervised with pseudo-labels $\hat{y}_{t}$ in the target domain or ground truth labels $y_{s}$ in the source domain. It is worth noting that the generated pseudo-labels can serve as supervision for the 1st-stage network as well, thus enabling domain adaptation for the 1st-stage network. Thus, the total loss of this network includes: 1. first-stage RoI regression loss $l_{reg1}$. 2. first-stage RoI classification loss $l_{cls1}$. 3. second-stage regression Smooth-L1 loss $l_{reg2}$. 4. second-stage classification loss $l_{cls2}$. 5. second-stage AU-NLL loss $l_{nll}$ specified in Sec IV-B.
III-2 Teacher Model
The teacher model processes raw (non-augmented) data and maintains fixed weights during the backward pass. Instead of employing standard backpropagation with predefined loss functions, the teacher model updates its weights using exponential moving average (EMA) from student model as follows:
@@ III-2 Teacher Model
Here, $\theta^{stu}_{t}$ represents the student model’s weights at iteration $t$, $\beta$ is the EMA decay factor that controls the update rate, and $\theta^{tea}_{t}$ denotes the teacher model’s weights. The resulting teacher model’s weights provide a smoothed representation of the student model’s weights over time.
@@ IV-A Enhancement of Pseudo-Label Quality
CTS starts by training a detector in a labeled source domain and then leverages this knowledge to generate high-quality pseudo-labels in the target domain. However, simulation-reality differences, such as differences in object size and point density, present significant challenges. Specifically, size bias has been shown to significantly reduce localization accuracy[13], especially when in simulation-to-reality scenarios, where expanding the simulation model asset library (e.g., through CAD modeling) to match the target domain is both difficult and expensive. To address these domain shifts and improve the reliability of pseudo-labels in the target domain, we propose the following methodology:
@@ IV-A1 Anchor Head (AH)
The second-stage model typically predicts size residuals $\Delta_{whl}$ between proposals from the first-stage and final bounding boxes, denoted as $\hat{B}$. This approach avoids regressing the size of bounding boxes entirely from scratch. However, a challenge arises when the first-stage model, trained with biased supervision from source domain labels, exhibits inaccuracy in estimating proposal sizes. Unreliable proposal box sizes can lead to size errors accumulating in the second stage, degrading final bounding box refinement accuracy and the effectiveness of pseudo-labels. Inspired by anchor-based detectors [24], we introduce a fixed-size anchor box ${w_{an},h_{an},l_{an}}$ to replace the proposal, termed the anchor head (AH). The AH replaces the traditional proposal mechanism, allowing the second-stage network to work with a globally fixed-size 3D anchor instead of refining proposals. By employing the AH in both the source and target domains, we ensure consistent behavior of the second-stage network regarding proposal size refinement across domains, thus avoid size error propagation in UDA and enhancing the quality of pseudo-labels.
@@ IV-A2 RoI Random Scaling (RRS) and Augmentation
where $\theta$ is the model parameter, $f_{\mu}$ and $f_{\sigma^{2}}$ represent subnetworks for predicting the mean and the variance.
@@
$\mathcal{L}_{nll}^{i}=\frac{\left(\overline{\mathbf{y}_{c}^{i}}-\overline{\hat{\mathbf{y}}_{c}^{i}}\right)^{2}}{2(\sigma_{c}^{i})^{2}}+\frac{1}{2}\log(\sigma_{c}^{i})^{2}$ (5)
Where final NLL loss $\mathcal{L}_{nll}=\frac{\sum_{i=1}^{8}\mathcal{L}_{nll}^{i}}{8}$ and uncertainty of box $u_{box}=\frac{\sum_{i=1}^{8}(\sigma_{c}^{i})^{2}}{8}$. And all components contribute equally to the loss and final uncertainty metric.
@@ IV-C Noise-aware Mean Teacher
Aligning transformations on both student-model inputs and teacher-model output facilitates the acquisition of domain-invariant representations, thereby aiding in adaptation to the target domain using pseudo-labels. However, noisy pseudo-labels can lead to error accumulation. To address this challenge, we leverage aleatoric uncertainties predicted by a model to annotate data in the target domain and mitigate the impact of noisy data during mean teacher domain adaptation with the following sampling strategies:
@@ IV-C2 Frame-Level Sampling
1:
2:$\mathcal{T}$: Unlabeled Target Domain Dataset
3:$\mathcal{T}_{sub}$: Target Sub-dataset after Sampling
4:$N_{t}$: Number of samples in $\mathcal{T}$
5:$N_{sub}$: Amount of data to be selected
6:$\mathcal{D}$: Noise-aware Model
7:while $N_{sub}<N_{t}$ do
8:     ${U}_{frame}\leftarrow\{\}$, $\mathcal{T}_{sub}\leftarrow\{\}$
9:     for each frame $x^{t}$ in $\mathcal{T}$ do
10:         $\hat{\mathbf{y}}^{t},\hat{\mathbf{u}}^{t}\leftarrow$ inference $\mathcal{D}$ for $x^{t}$
11:         $\hat{u}^{t}\leftarrow\text{mean of }\hat{\mathbf{u}}^{t}\text{ for all valid object in $x^{t}$}$
12:         ${U}_{frame}\leftarrow$ append $\hat{u}_{t}$ to ${U}_{frame}$
13:     end for
14:     for $i$ in $\{1,\ldots,{N}_{sub}\}$ do
15:         $j\leftarrow\text{argmin of }{U}_{frame}$
16:         $\mathcal{T}_{sub}\leftarrow$ append the $j_{th}$ element $x_{j}^{t}$ in $\mathcal{T}$ to $\mathcal{T}_{sub}$
17:         pop the $j_{th}$ element $\hat{u}_{j}^{t}$ from ${U}_{frame}$
18:     end for
19:     $\mathcal{D}\leftarrow$ fine-tune $\mathcal{D}$ with $\mathcal{T}_{sub}$
20:     ${N}_{sub}+={N}_{sub}$
21:end while
22:return $\mathcal{D}$
@@ V-A1 Datasets
Most existing LiDAR simulation datasets are primarily used for task-specific problems, such as Vehicle-to-Vehicle Communication [28] and Continuous Domain Shift [29], rather than for sim-to-real UDA as addressed in this paper. Thus, we conduct supervised training in a simulated source domain, namely CARLA3D, acquired within the CARLA simulator [9] from scratch. All samples are taken from eight built-in scenarios in CARLA to ensure data diversity. The ego-vehicle is positioned randomly, collecting about 100 samples per scenario, each comprising eight frames at 2Hz. Out of the eight frames per sample, five are randomly chosen for the training set, yielding 3,990 frames with a total of 25,192 objects. Further details of the CARLA3D dataset are outlined in Table II. The target domains chosen include KITTI [10], Lyft [11], and TinySUScape used in [12]. During the testing phase, samples from these datasets along with their corresponding labels will be utilized, whereas only samples will be used during the training phase. A summary of these datasets is presented in Table III.
@@ V-B Main Results
Our CTS framework was compared with the following methods: 1. SN[13]: A domain adaptation method has been considered effective on various datasets; 2. MLC-Net[6]: A domain adaptation method also based on mean teacher, which is similar to ours in the mean teacher part; 3. ST3D++[8]: A recent self-training based method that achieved state-of-the-art performance in real-to-real (e.g., Nuscenes[30] $\to$ KITTI[10]) domain adaptation tasks.
@@ V-C4 Benefits of Noise Awareness in Mean Teacher
Figure 6: Examples of different levels of difficulties in 3D boxes. The blue boxes represent the ground truth; the green boxes represent the predicted results. The points in different colors at the box corners represent the 8 AU value components, whose mean is the final AU value of the entire object.
@@ VI Conclusion
This paper has introduced a CTS framework for unsupervised domain adaptation (UDA) in 3D object detection, bridging the gap between simulation and real-world domains. The proposed techniques, including RoI random scaling and augmentation, along with the fixed-size anchor head, enhance the diversity of simulation data and address object size discrepancies across domains, thereby improving the quality of pseudo-labels. Additionally, the proposed aleatoric uncertainty (AU) estimation, based on a uniform corner-format representation of bounding boxes, facilitates the integration of pseudo-label noise awareness into the mean teacher domain adaptation process, leading to high-quality pseudo-label sampling. Experimental results on the CARLA, KITTI, Lyft, and TinySUScape datasets demonstrate substantial improvements over existing methods in various sim-to-real UDA tasks, with 5%-17% gains in $AP_{3D}$ and 2%-10% gains in $AP_{BEV}$. Future work will focus on extending this approach to cover both sim-to-real and real-to-real UDA scenarios, as well as incorporating additional categories (e.g., bicycles, pedestrians) in domain adaptation.